OpenStudy (anonymous):

Calculus Help: Show that ln(y)=arctan(x) is a solution to the differential equation (1+x^2)y"+(2x-1)y'=0 I'm totally lost need help!

OpenStudy (mathmate):

From ln(y)=arcxtan(x), raise to power of e e^(ln(y))=e^(arctan(x)) y=e^arctan(x) now differentiate y to get y', and differentiate twice to get y". Substitute in the given equation's LHS to show that after simplification, the LHS = 0 (for any value of x).

OpenStudy (mathmate):

Hint: you will need the chain rule to do the differentiations.

OpenStudy (anonymous):

Thank you so much mathmate! you're a lifesaver :)

OpenStudy (mathmate):

You're welcome! :)